# Perceptions of the Fairness Impacts of Multiplicity in Machine Learning

### v1

@@ Abstract
Machine learning (ML) is increasingly used in high-stakes settings, yet multiplicity – the existence of multiple good models – means that some predictions are essentially arbitrary. ML researchers and philosophers posit that multiplicity poses a fairness risk, but no studies have investigated whether stakeholders agree. In this work, we conduct a survey to see how the presence of multiplicity impacts lay stakeholders’ – i.e., decision subjects’ – perceptions of ML fairness, and which approaches to address multiplicity they prefer. We investigate how these perceptions are modulated by task characteristics (e.g., stakes and uncertainty). Survey respondents think that multiplicity lowers distributional, but not procedural, fairness, even though existing work suggests the opposite. Participants are strongly against resolving multiplicity by using a single good model (effectively ignoring multiplicity) or by randomizing over possible outcomes. Our results indicate that model developers should be intentional about dealing with multiplicity in order to maintain fairness.
1 Introduction
@@
Machine learning is ubiquitous in high-stakes decision-making settings, e.g., social services [18], criminal justice [75], healthcare [71], and hiring [59]. Due to the life-altering impacts of these decisions (for instance, whether someone can get hired at a job), it is important that the model making the decision is as accurate, fair, and accountable as possible. The conventional machine learning wisdom is that there exists a single best-performing model and that achieving other desiderata like fairness or robustness requires an accuracy tradeoff [67]. But the literature – both recent [25, 65] and foundational [11] – challenges that paradigm: there are typically many models that achieve similar predictive accuracy by different means (i.e., by making mistakes on different parts of the input distribution). This ambiguity leads to predictive multiplicity, which is when two equally good models assign different predictions to a given sample, as illustrated by fig. 1. In socially-salient prediction settings, the set of (approximately) equally good models is often very large [65, 77, 100], and a large fraction of samples receive conflicting outcomes under this multiplicity [65].
@@
Work in computer science and philosophy suggests that multiplicity poses a fairness risk [8, 22, 23]. Fairness has multiple components, including distributional fairness, which considers whether the outcomes are fair, and procedural fairness, which considers whether the decision was made in a fair way, regardless of outcomes [20]. The existing work points out that multiplicity can violate procedural fairness, because choosing one model over the alternatives results in arbitrary decisions [8, 22, 23]. This fairness violation is especially salient when models operate in high-stakes settings where the justifiability of a negative decision is important. For instance, suppose a company uses an ML model to determine which job applicants to interview, but that there is a different model that would do equally as well at selecting promising candidates (e.g., and have similar accuracy on the available validation data). A candidate who is not interviewed according to , but who would have been interviewed if is used, may feel that procedural fairness is violated if the company did not even consider using . An additional concern is that when models are deployed at scale, multiplicity can result in systematic arbitrariness that locks people out of opportunities [23]. For instance, if all companies contract with the same software vendor and each uses model to determine who to interview, may get no job interviews even though the existence of suggests they may be qualified.
@@
While the existing literature makes a strong argument for the fairness risks of multiplicity, no research has interrogated whether lay stakeholders believe that multiplicity impacts the fairness of machine learning models. Likewise, no research has asked how stakeholders think decisions should be made when a machine learning system contains multiplicity. More precisely, consider the same job application scenario where two models and have similar performance, yet disagree on whether to interview a job candidate . The standard machine learning approach is to choose a single well-performing model, for instance, selecting without even realizing that exists. Philosophers suggest using randomization – e.g., flipping a coin to determine whether to interview – as the fairest way to deal with multiplicity [23, 44]. Computer scientists suggest switching to more sophisticated model architectures to reduce multiplicity (e.g., ensemble methods as a way to reduce variance) [7, 22, 76]. Lay stakeholders will not necessarily like those approaches, as research shows that different types of stakeholders (like decision makers and decision subjects) can have different perceptions of what matters for fairness in ML [57, 84, 102]. Critical scholars argue that considering stakeholder preferences is important [10, 87] and this belief underlies frameworks like participatory ML [31] and value-sensitive design [94]. Furthermore, companies deploying ML may care about public opinions of fairness to maintain high levels of trust in their product [99].
@@
To address these questions, we want to evaluate fairness perceptions for a variety of tasks. However, some of the factors we care about for RQ3 (namely, task stakes and uncertainty) may be subjective, so we did not want to rely solely on our own classification of tasks. Instead, we conducted a preliminary study with participants, which, to the best of our knowledge, is the first study to get lay stakeholders’ opinions on both tasks’ perceived stakes and uncertainty (Castelo et al. [15] collect opinions on task uncertainty, but not stakes). Based on the results of this survey and our own assessment of task framing (reward vs. punishment), we selected 8 diverse tasks to use in the main study.
@@
Our results indicate that stakeholders do not perceive multiplicity as an overall fairness risk, counter to what the literature suggests. However, this result is nuanced as participants are more likely to see distributional, rather than procedural, fairness harms as a consequence of multiplicity. We also find that participants have clear preferences for how multiplicity is resolved namely by deferring to a human expert for situations where multiplicity impacts the predictions (RQ2). These preferences vary depending on the specific task, with task stakes and task framing significantly impacting resolution preferences (RQ3). Despite philosophical arguments for randomization as a multiplicity solution [23, 44], this was the least preferred option in our experiment.
These results have two important implications: first, the way that multiplicity is discussed in the existing literature does not align with the lay stakeholders’ perceptions. Instead, we hypothesize that other concerns about machine learning (such as recourse and explainability) may dominate participants’ fairness assessments. Second, even though overall fairness perceptions do not shift significantly, the strong sentiments that participants have regarding multiplicity resolution suggest that ML developers must be deliberate in determining how models behave in the presence of multiplicity.
2 Background of Fairness and Multiplicity
@@
2.1 Fairness in machine learning
Colquitt and Rodell propose several components of fairness including distributional, procedural, and informational fairness [21]. Distributional fairness asks whether a model’s outcomes are fair, irrespective of how the decisions were made. For instance, group fairness metrics such as equalized odds [29, 38] ask whether different demographic groups are treated equally in aggregate. Procedural fairness, by contrast, asks whether a decision was made in a fair way, regardless of what the final outcome is. Procedural fairness has many facets, including counterfactual fairness [53] (how a test sample needs to change to get a different prediction) and the choice of features that are included in a model. For example, in a task such as hiring, procedural fairness may be violated if race is a decisive factor (i.e., changing an applicant’s race changes whether they are hired) or even if race is included as a model input at all. Informational fairness requires that details about the decision making process be available, and has been explored in ML through the use of explanations, e.g. [6, 28, 83, 101]. In this paper, we consider perceptions of distributional and procedural fairness by asking whether outcomes are fair and whether the use of the ML model is appropriate.
2.2 Multiplicity in machine learning and fairness concerns
Predictive multiplicity [65], which we simply call multiplicity for brevity, occurs when there are multiple ML models that are equally well-suited for a task, yet give different individual-level predictions. Formally, suppose we have a dataset , a hypothesis class (e.g., all possible decision trees with depth ) and some loss function (e.g., the 0-1 loss). A standard machine learning pipeline aims to find that minimizes (e.g., the total number of misclassifications on ). This solution is typically found via an approximate optimization process such as gradient descent, and yields a model that hopefully has close-to-optimal loss, i.e., has . But sometimes, the optimization problem is underconstrained and there is another model such that . (Intuitively, there may be multiple global minima.) If there is some input with we say that the models exhibit (predictive) multiplicity.
@@ 2.2 Multiplicity in machine learning and fairness concerns
Computing the fraction of inputs whose predictions are affected by multiplicity is generally intractable, so existing approaches quantify multiplicity by limiting their analysis to specific model types, approximating the impacts of multiplicity, and/or using brute force. Estimates on the COMPAS dataset, which is used to predict risk of recidivism in defendants, indicate that multiplicity affects 44% of test inputs when using linear models [65] and 69.4% of test inputs when using decision trees [22]. For probabilistic models that compute risk scores, model outputs vary by on 82-100% of samples across three tabular datasets [97]. These levels of multiplicity indicate that many predictions are a function of the specific model that is chosen, rather than purely a function of the data.
@@
Fairness risks of multiplicity include harms to an individual, society or a group of individuals, and the entity who is deploying the model. Fairness risks to individuals include a lack of justifiability for a negative decision [8] and systemic loss of opportunity when models containing multiplicity are deployed at scale [23]. Society or groups of individuals may be adversely impacted by multiplicity if a sub-optimal model, with regards to fairness, is selected from the set of equivalently-accurate models. The entity using the model is also at risk, first for potential lawsuits or reputational harm if their decisions are not justifiable [8]. Second, if a single model is deployed at scale when multiplicity exists, this standardization can provide a worse set of outcomes, e.g., companies all using the same recruiting software may end up with lower-quality new hires than if each company used its own model [52]. However, no work that we are aware of has investigated the fairness risks of multiplicity from the perspective of stakeholders. Previous work shows that formal fairness definitions often do not line up with stakeholders’ fairness perceptions [90], so, it seems likely that experts and lay stakeholders may have distinct insights about whether multiplicity is (un)fair.
2.3 “Solving” multiplicity
@@
Refining the modeling technique can reduce multiplicity, such as by using cascading classifiers [76], bootstrapping [22], or selective ensembles that abstain if the confidence is not high enough [7]. Similar ensembling approaches have long been used to improve accuracy in ML applications [78]. The difference when using ensembling to combat multiplicity is that multiplicity is often discussed in the context of social prediction tasks where an objective ground truth is elusive or even does not exist. Thus, using ensembling to get rid of multiplicity often does not yield significantly higher accuracy than using a non-ensembled model. It does, however, mask that multiplicity exists, making it harder to challenge the validity of an algorithm. We refer to ensembling and other approaches that aim to reduce multiplicity through modeling choices as fancy, i.e., training a fancier (more complex) model.
@@
ML researchers have suggested abstaining, and, presumably, deferring to a human expert (denoted in this paper as human), when a decision is subject to multiplicity [7]. This approach parallels a long line of work in robust machine learning that similarly suggests abstaining on decisions when robustness cannot be verified [3, 19, 40, 55].222Note that non-robustness constitutes a form of multiplicity, as it implies there is another reasonable model that outputs an alternate prediction for the test sample. Within the ML literature, robustness is typically discussed for deep networks and/or text and image data, while multiplicity research largely focuses on social prediction tasks using tabular data. However, these works rarely present a plan for what abstaining looks like in practice: often, abstaining on a decision results in not taking action, thereby defaulting to the negative label (for instance, abstaining on a loan application is effectively denying it). We consider abstaining and deferring to a default option in our survey. However, attitudes may vary depending on whether the outcome is desirable or not, so we use default-good and default-bad to represent the outcomes that are good and bad, respectively, from the perspective of the decision subject.
An alternate procedural approach is random, i.e., randomizing over the set of equally-good models [23, 44]. ML in social prediction tasks often serves to distribute a limited resource, e.g., admissions to a college with a fixed enrollment limit. Lotteries are typically viewed as fair system for dividing limited resources [12, 88]. However, there are technical challenges in implementing randomization effectively [44].
@@
Traditional ML practice assumes that there is a trade-off between model accuracy and fairness, as defined by group-level metrics such as equalized odds or statistical parity [47, 67]. Recent research to explore the whole set of equally-good models challenges this paradigm, showing that accuracy and other criteria are not in tension [77, 100]. But even if we can use multiplicity to find fairer (or more robust) models, there are still multiple mutually incompatible fairness (or robustness) definitions. We do not include optimizing for fairness as one of the possible multiplicity resolution techniques because the task of choosing the fairest model is typically not well-defined [17, 51].
3 Related Work and Hypotheses
Fairness perceptions of machine learning algorithms have been studied extensively, as evidenced by multiple survey papers on the topic [68, 91]. Here, we will focus on dimensions most related to our research questions.
Research in this area investigates distributional fairness (e.g., matching lay stakeholder’s fairness notions to formal definitions [82, 90]), procedural fairness (e.g., considering what features to include [35]), and informational fairness (e.g., via explanations [83, 101]). In this work, we focus on perceptions of procedural and distributional fairness, presenting multiplicity as a hurdle that must be dealt with procedurally but that also affects model outcomes. The fairness of an algorithm is often considered relative to a baseline of human decision making. Fairness perceptions of a model relative to a human decision maker can either increase from algorithmic appreciation [60], or decrease from algorithmic aversion [27]. Existing work establishes that fairness perceptions of ML vary based on task characteristics, including stakes [37, 81, 101], uncertainty [80], complexity [79], and framing (reward vs. punishment) [37].

### v2

@@ Abstract.
Machine learning (ML) is increasingly used in high-stakes settings, yet multiplicity – the existence of multiple good models – means that some predictions are essentially arbitrary. ML researchers and philosophers posit that multiplicity poses a fairness risk, but no studies have investigated whether stakeholders agree. In this work, we conduct a survey to see how multiplicity impacts lay stakeholders’ – i.e., decision subjects’ – perceptions of ML fairness, and which approaches to address multiplicity they prefer. We investigate how these perceptions are modulated by task characteristics (e.g., stakes and uncertainty). Survey respondents think that multiplicity threatens the fairness of model outcomes, but not the appropriateness of using the model, even though existing work suggests the opposite. Participants are strongly against resolving multiplicity by using a single model (effectively ignoring multiplicity) or by randomizing the outcomes. Our results indicate that model developers should be intentional about dealing with multiplicity in order to maintain fairness.
1. Introduction
@@
Machine learning (ML) is ubiquitous in high-stakes decision-making settings, e.g., social services (Chouldechova et al., 2018), criminal justice (Propublica, 2016), healthcare (Obermeyer et al., 2019), and hiring (Li et al., 2021). Due to the life-altering impacts of these decisions (for instance, whether someone can get hired at a job), it is important that the model making the decision is as accurate, fair, and accountable as possible. The conventional machine learning wisdom is that there exists a single best-performing model and that achieving other desiderata like fairness or robustness requires an accuracy tradeoff (Menon and Williamson, 2018). But the literature – both recent (D’Amour et al., 2022; Marx et al., 2020) and foundational (Breiman, 2001) – challenges that paradigm: there are typically many models that achieve similar predictive accuracy by different means (i.e., by making mistakes on different parts of the input distribution). This ambiguity leads to predictive multiplicity, which is when two equally good models assign different predictions to a given sample, as illustrated by fig. 1. In socially-salient prediction settings, the set of (approximately) equally good models is often very large (Marx et al., 2020; Rudin et al., 2024; Xin et al., 2022), and a large fraction of samples receive conflicting outcomes under this multiplicity (Marx et al., 2020).
@@
Work in computer science and philosophy suggests that multiplicity poses a fairness risk (Black et al., 2022b; Cooper et al., 2024; Creel and Hellman, 2022). Fairness has multiple components, including distributional fairness, which considers whether the outcomes are fair, and procedural fairness, which considers whether the decision was made in a fair way, regardless of outcomes (Colquitt, 2001). The existing work points out that multiplicity can violate procedural fairness, because choosing one model over the alternatives results in arbitrary decisions (Black et al., 2022b; Cooper et al., 2024; Creel and Hellman, 2022). This fairness violation is especially salient when models operate in high-stakes settings where the justifiability of a negative decision is important. For instance, suppose a company uses an ML model to determine which job applicants to interview, but that there is a different model that would do equally as well at selecting promising candidates (e.g., and have similar accuracy on the available validation data). A candidate who is not interviewed according to , but who would have been interviewed if were used, may feel that procedural fairness is violated if the company did not even consider using . An additional concern is that when models are deployed at scale, multiplicity can result in systematic arbitrariness that locks people out of opportunities (Creel and Hellman, 2022). For instance, if all companies contract with the same software vendor and each uses model to determine who to interview, may get no job interviews even though the existence of suggests they may be qualified.
@@
While the existing literature makes a strong argument for the fairness risks of multiplicity, no research has interrogated whether lay stakeholders believe that multiplicity impacts the fairness of machine learning models. Likewise, no research has asked how stakeholders think decisions should be made when a machine learning system contains multiplicity. More precisely, consider the same job application scenario where two models and have similar performance, yet disagree on whether to interview a job candidate . The standard machine learning approach is to choose a single well-performing model, for instance, selecting without even realizing that exists. Philosophers suggest using randomization – e.g., flipping a coin to determine whether to interview – as the fairest way to deal with multiplicity (Creel and Hellman, 2022; Jain et al., 2024). Computer scientists suggest switching to more sophisticated model architectures to reduce multiplicity (e.g., ensemble methods as a way to reduce variance) (Black et al., 2022a; Cooper et al., 2024; Roth et al., 2023). Lay stakeholders will not necessarily like those approaches, as research shows that different types of stakeholders (like decision makers and decision subjects) can have different perceptions of what matters for fairness in ML (Lee and Rich, 2021; Scott et al., 2021; Zhang et al., 2023). Critical scholars argue that considering stakeholder preferences is important (Bondi et al., 2021; Selbst et al., 2019) and this belief underlies frameworks like participatory ML (Feffer et al., 2023) and value-sensitive design (Umbrello and Van de Poel, 2021). Furthermore, companies deploying ML may care about public opinions of fairness to maintain high levels of trust in their product (Woodruff et al., 2018).
@@
The image contains 3 sub-figures, labeled A, B, and C. Image A has a cluster of data points (red circles and blue triangles) that are not perfectly linearly separable. The image shows two linear classifiers that each misclassify 1 point (but a different point for each). The points that are misclassified by a model are circled in yellow. Image B is formatted as a table where headers are pictures of animals (Column 1 Cat, Column 2 Dog, Column 3 cat). Each row contains the prediction that an AI system makes on each image (predicting whether it is a cat or dog). The first row (AI 1) is correct on examples 1 and 3, while the second row (AI 2) is correct on examples 2 and 3. Text in the image indicates that both AIs are correct on 2 out of 3 examples. Image C is a venn diagram. There is a large circle labeled All Reviews. Within this circle, there are 3 smaller circles arranged along an invisible diagonal line. The middle circle (green) is labeled Reviews Deleted by Human Moderator. The other circles are labeled Reviews deleted by the original model (orange) and Reviews deleted by the new model (pink). The orange and pink circles overlap about 50 percent of their area, while each overlaps with the green circle by about 75 percent of their area.
@@
To address these questions, we want to evaluate fairness perceptions for a variety of tasks. However, some of the factors we care about for RQ3 (namely, task stakes and uncertainty) may be subjective, so we did not want to rely solely on our own classification of tasks. Instead, we conducted a preliminary study with participants, which, to the best of our knowledge, is the first study to get lay stakeholders’ opinions on both tasks’ perceived stakes and uncertainty (Castelo et al. (Castelo et al., 2019) collect opinions on task uncertainty, but not stakes). Based on the results of this survey and our own assessment of task framing (reward vs. punishment), we selected 8 diverse tasks to use in the main study.
@@
Our results indicate that stakeholders do not perceive multiplicity as an overall fairness risk, counter to what the literature suggests. However, this result is nuanced as participants are more likely to express that the model outcomes might be unfair, rather than that the model may be inappropriate to use at all. We also find that participants have clear preferences for how multiplicity is resolved, namely by deferring to a human expert for situations where multiplicity impacts the predictions (RQ2). These preferences vary depending on the specific task, with task stakes and task framing significantly impacting resolution preferences (RQ3). Despite philosophical arguments for randomization as a multiplicity solution (Creel and Hellman, 2022; Jain et al., 2024), this was the least preferred option in our experiment.
These results have two important implications: first, the way that multiplicity is discussed in the existing literature does not align with the lay stakeholders’ perceptions. Instead, we hypothesize that other concerns about machine learning (such as recourse and explainability) may dominate participants’ fairness assessments. Second, even though overall fairness perceptions do not shift significantly after learning about multiplicity, the strong sentiments that participants have regarding multiplicity resolution suggest that ML developers must be deliberate in determining how models behave in the presence of multiplicity. In particular, simply choosing a single well-performing model and not exploring the impacts of multiplicity is viewed as unfair.
2. Background of Fairness and Multiplicity
@@
2.1. Fairness in machine learning
Colquitt and Rodell propose several components of fairness including distributional, procedural, and informational fairness (Colquitt and Rodell, 2015). Distributional fairness asks whether a model’s outcomes are fair, irrespective of how the decisions were made. For instance, group fairness metrics such as equalized odds (Dwork et al., 2012; Hardt et al., 2016) ask whether different demographic groups are treated equally in aggregate. Procedural fairness, by contrast, asks whether a decision was made in a fair way, regardless of what the final outcome is. Procedural fairness has many facets, including counterfactual fairness (Kusner et al., 2017) (how a test sample needs to change to get a different prediction) and the choice of features that are included in a model. For example, in a task such as hiring, procedural fairness may be violated if race is a decisive factor (i.e., changing an applicant’s race changes whether they are hired) or even if race is included as a model input at all. Informational fairness requires that details about the decision making process be available, and has been explored in ML through the use of explanations, e.g. (Binns et al., 2018; Dodge et al., 2019; Schoeffer et al., 2022; Yurrita et al., 2023). In this paper, we consider perceptions of procedural fairness.
The fairness of a ML model can be measured based on its adherence to some fixed standard (e.g., a group fairness metric), or in a human-centered way through perceived fairness, as in this study. Perceived fairness is a common way to evaluate ML systems (Starke et al., 2022), perhaps because settling on a formal fairness definition is challenging. Research has established that perceived fairness of a model correlates with trust of that model (Shin et al., 2020; Shin and Park, 2019; Woodruff et al., 2018), and suggests that low perceived fairness corresponds to negative consequences (e.g., reputational harm) for the entity deploying ML (Starke et al., 2022).
2.2. Multiplicity in machine learning and fairness concerns
Predictive multiplicity (Marx et al., 2020), which we simply call multiplicity for brevity, occurs when there are multiple ML models that are equally well-suited for a task, yet give different individual-level predictions. Formally, suppose we have a dataset , a hypothesis class (e.g., all possible decision trees with depth ) and some loss function (e.g., the 0-1 loss). A standard machine learning pipeline aims to find that minimizes (e.g., the total number of misclassifications on ). This solution is typically found via an approximate optimization process such as gradient descent, and yields a model that hopefully has close-to-optimal loss, i.e., has . But sometimes, the optimization problem is underconstrained and there is another model such that . (Intuitively, there may be multiple global minima.) If there is some input with we say that the models exhibit (predictive) multiplicity.
@@ 2.2. Multiplicity in machine learning and fairness concerns
Computing the fraction of inputs whose predictions are affected by multiplicity is generally intractable, so existing approaches quantify multiplicity by limiting their analysis to specific model types, approximating the impacts of multiplicity, and/or using brute force. Estimates on the COMPAS dataset, which is used to predict risk of recidivism in defendants, indicate that multiplicity affects 44% of test inputs when using linear models (Marx et al., 2020) and 69.4% of test inputs when using decision trees (Cooper et al., 2024). For probabilistic models that compute risk scores, model outputs vary by on 82-100% of samples across three tabular datasets (Watson-Daniels et al., 2023). These levels of multiplicity indicate that many predictions are a function of the specific model that is chosen, rather than purely a function of the data.
@@
Fairness risks of multiplicity include harms to an individual, society or a group of individuals, and the entity who is deploying the model. Fairness risks to individuals include a lack of justifiability for a negative decision (Black et al., 2022b) and systemic loss of opportunity when models containing multiplicity are deployed at scale (Creel and Hellman, 2022).
Example 2.1 (Lack of justifiability).
Black et al (Black et al., 2022b) point out that US credit laws like the Equal Credit Opportunity Act aim to reduce arbitrariness and increase equity in credit decisions, such as by requiring lenders to provide justification for denied loan requests. However, if another multiple equally-accurate model approves a loan request, it may be more challenging to justify why the bank needed to use the model that denies the loan.
Example 2.2 (Systemic loss of opportunity).
Creel and Hellman (Creel and Hellman, 2022) discuss how individual employers may have idiosyncratic hiring criteria, but that the harm from this to job applicants is minimal, as they can apply to many jobs. Since each employer has their own potentially arbitrary criteria (in other words, multiplicity exists in the process of choosing who to hire), the job applicant will eventually be hired by someone. However, if every company uses the same hiring system, then certain job applicants will not be able to be hired anywhere, causing them significant harm. This setup can happen, for instance, if every company contracts with the same third-party software to screen resumes.
Society or groups of individuals may be adversely impacted by multiplicity if a sub-optimal model, with regards to fairness, is selected from the set of equivalently-accurate models. The entity using the model is also at risk, first for potential lawsuits or reputational harm if their decisions are not justifiable (Black et al., 2022b). Second, if a single model is deployed at scale when multiplicity exists, this standardization can provide a worse set of outcomes.
Example 2.3 (Worse outcomes for the decision-maker).
Kleinberg et al. (Kleinberg and Raghavan, 2021) show that companies who all use the same recruiting software may end up with lower-quality new hires than if each company used its own model, an outcome that is harmful for the company itself and not just potential job applicants.
However, no work that we are aware of has investigated the fairness risks of multiplicity from the perspective of stakeholders. Previous work shows that formal fairness definitions often do not line up with stakeholders’ fairness perceptions (Srivastava et al., 2019), so, it seems likely that experts and lay stakeholders may have distinct insights about whether multiplicity is (un)fair.
2.3. “Solving” multiplicity
@@
Refining the modeling technique can reduce multiplicity, such as by using cascading classifiers (Roth et al., 2023), bootstrapping (Cooper et al., 2024), or selective ensembles that abstain if the confidence is not high enough (Black et al., 2022a). Similar ensembling approaches have long been used to improve accuracy in ML applications (Sagi and Rokach, 2018). The difference when using ensembling to combat multiplicity is that multiplicity is often discussed in the context of social prediction tasks where an objective ground truth is elusive or even does not exist. Thus, using ensembling to get rid of multiplicity often does not yield significantly higher accuracy than using a non-ensembled model. It does, however, mask that multiplicity exists, making it harder to challenge the validity of an algorithm. We refer to ensembling and other approaches that aim to reduce multiplicity through modeling choices as fancy, i.e., training a fancier (more complex) model.
@@
ML researchers have suggested abstaining, and, presumably, deferring to a human expert (denoted in this paper as human), when a decision is subject to multiplicity (Black et al., 2022a). However, the design of “human-in-the-loop systems” is not straightforward (Wu et al., 2022), and may interact with other criteria we are trying to measure, e.g., algorithm aversion (Fogliato et al., 2022). The idea of abstaining and deferring to a human decision-maker approach parallels a long line of work in robust machine learning that similarly suggests abstaining on decisions when robustness cannot be verified (Bartlett and Wegkamp, 2008; Cohen et al., 2019; Herbei and Wegkamp, 2006; Laidlaw and Feizi, 2019).222Note that non-robustness constitutes a form of multiplicity, as it implies there is another reasonable model that outputs an alternate prediction for the test sample. Within the ML literature, robustness is typically discussed for deep networks and/or text and image data, while multiplicity research largely focuses on social prediction tasks using tabular data. However, these works rarely present a plan for what abstaining looks like in practice: often, abstaining on a decision results in not taking action, thereby defaulting to the negative label (for instance, abstaining on a loan application is effectively denying it). We consider abstaining and deferring to a default option in our survey. However, attitudes may vary depending on whether the outcome is desirable or not, so we use default-good and default-bad to represent the outcomes that are good and bad, respectively, from the perspective of the decision subject.
An alternate procedural approach is random, i.e., randomizing over the set of equally-good models (Creel and Hellman, 2022; Jain et al., 2024). ML in social prediction tasks often serves to distribute a limited resource, e.g., admissions to a college with a fixed enrollment limit. Lotteries are typically viewed as fair system for dividing limited resources (Broome, 1990; Sher, 1980). However, there are technical challenges in implementing randomization effectively (Jain et al., 2024).
@@
Traditional ML practice assumes that there is a trade-off between model accuracy and fairness, as defined by group-level metrics such as equalized odds or statistical parity (Kamiran and Calders, 2012; Menon and Williamson, 2018). Recent research to explore the whole set of equally-good models challenges this paradigm, showing that accuracy and other criteria are not in tension (Rudin et al., 2024; Xin et al., 2022). But even if we can use multiplicity to find fairer (or more robust) models, there are still multiple mutually incompatible fairness (or robustness) definitions. We do not include optimizing for fairness as one of the possible multiplicity resolution techniques because the task of choosing the fairest model is typically not well-defined (Chouldechova, 2017; Kleinberg et al., 2017).
3. Related Work and Hypotheses
Fairness perceptions of machine learning algorithms have been studied extensively, as evidenced by multiple survey papers on the topic (Narayanan et al., 2024; Starke et al., 2022). Here, we will focus on dimensions most related to our research questions.
Research in this area investigates distributional fairness (e.g., matching lay stakeholder’s fairness notions to formal definitions (Saxena et al., 2019; Srivastava et al., 2019)), procedural fairness (e.g., considering what features to include (Grgic-Hlaca et al., 2018)), and informational fairness (e.g., the impact of explanations (Schoeffer et al., 2022; Yurrita et al., 2023)). In this work, we focus on perceptions of procedural fairness, presenting multiplicity as a hurdle that must be dealt with procedurally. The fairness perceptions of ML models are often considered relative to a baseline of human decision making, and can either increase from algorithmic appreciation (Logg et al., 2019), or decrease from algorithmic aversion (Dietvorst et al., 2015). Existing work establishes that fairness perceptions of ML models vary based on task characteristics, including stakes (Hannan et al., 2021; Saragih and Morrison, 2022; Yurrita et al., 2023), uncertainty (Salimzadeh et al., 2024), complexity (Salimzadeh et al., 2023), and framing (reward vs. punishment) (Hannan et al., 2021).